Q) A train is travelling between two stations that are 100km apart at a speed of V kmh^-1.
(1) express the time taken for the journey in terms of V

Answer by 303795(602)

Time is distance divided by speed
ie Time = Distance / Speed
= 100 /V with the answer measured in hours.
eg if speed is 100 kmh^-1 then time = 100/100 = 1 hour.
If speed = 50 kmh^-1 then time = 100/50 = 2 hours.
If speed = 2 kmh^-1 then time = 100/2 = 50 hours.